Suppose A and B are 3x3 matrices. Prove that (i) (A+B)^T = B^T +A^T
(ii) det (AB)^T= det(A).det(B)
1
Expert's answer
2020-10-07T18:38:48-0400

(i) Let "A=\left(\begin{array}{ccc} a_{11} & a_{12} & a_{13} \\ a_{21} & a_{22} & a_{23} \\ a_{31} & a_{32} & a_{33} \end{array}\right), \ \ \ 
B=\left(\begin{array}{ccc} b_{11} & b_{12} & a_{13} \\ b_{21} & b_{22} & b_{23} \\ b_{31} & b_{32} & b_{33} \end{array}\right)."


Then "(A+B)^T=\left(\begin{array}{ccc} a_{11}+b_{11} & a_{12}+b_{12} & a_{13}+b_{13} \\ a_{21}+b_{21} & a_{22}+b_{22} & a_{23}+b_{23} \\ a_{31}+b_{31} & a_{32}+b_{32} & a_{33}+b_{33} \end{array}\right)^T ="


"=\left(\begin{array}{ccc} a_{11}+b_{11} & a_{21}+b_{21} & a_{31}+b_{31} \\ a_{12}+b_{12} & a_{22}+b_{22} & a_{32}+b_{32} \\ a_{13}+b_{13} & a_{23}+b_{32} & a_{33}+b_{33} \end{array}\right)="


"=\left(\begin{array}{ccc} b_{11}+a_{11} & b_{21}+a_{21} & b_{31}+a_{31} \\ a_{12}+b_{12} & b_{22}+ a_{22} & b_{32}+a_{32} \\ b_{13}+a_{13} & b_{32}+a_{23} & b_{33}+ a_{33} \end{array}\right)="


"=\left(\begin{array}{ccc} b_{11} & b_{21} & b_{31} \\ b_{12} & b_{22} & b_{32} \\ b_{13} & b_{32} & b_{33} \end{array}\right)+\left(\begin{array}{ccc} a_{11} & a_{21} & a_{31} \\ a_{12} & a_{22} & a_{32} \\ a_{13} & a_{23} & a_{33} \end{array}\right)=B^T+A^T."



(ii) Taking into account that "\det(M^T)=\det (M)" and "\det(MN)=\det(M)\det(N)" for any "3\times 3" matrix "M" and "N", we conclude that "\det (AB)^T=\det(AB)= \det(A)\det(B)."